OpenStudy (anonymous):

How would I use the Pythagorean Identities to simplify: cos squared theta + tan squared theta cos squared theta and cos squared theta divided by 1-sin theta

OpenStudy (dan815):

use latex or draw your question

OpenStudy (bibby):

\[\cos^2(\theta) + \tan^2(\theta)\cos^2(\theta)\] this alone is one problem, right?

OpenStudy (anonymous):

Yes

OpenStudy (dan815):